# Emergent Mixture-of-Experts: Can Dense Pre-trained Transformers Benefit from Emergent Modular Structures?

### v1

@@ Abstract
Incorporating modular designs into neural networks demonstrates superior out-of-generalization, learning efficiency, etc. Existing modular neural networks are generally explicit because their modular architectures are pre-defined, and individual modules are expected to implement distinct functions. Conversely, recent works reveal that there exist implicit modular structures in standard pre-trained transformers, namely Emergent Modularity. They indicate that such modular structures exhibit during the early pre-training phase and are totally spontaneous. However, most transformers are still treated as monolithic models with their modular natures underutilized. Therefore, given the excellent properties of explicit modular architecture, we explore whether and how dense pre-trained transformers can benefit from emergent modular structures. To study this question, we construct Emergent Mixture-of-Experts (EMoE). Without introducing additional parameters, EMoE can be seen as the modular counterpart of the original model and can be effortlessly incorporated into downstream tuning. Extensive experiments (we tune 1785 models) on various downstream tasks (vision and language) and models (22M to1.5B) demonstrate that EMoE effectively boosts in-domain and out-of-domain generalization abilities. Further analysis and ablation study suggest that EMoE mitigates negative knowledge transfer and is robust to various configurations. Code is available at https://github.com/qiuzh20/EMoE
@@ 1 Introduction
Modularity attracts considerable attention from the artificial intelligence community (Auda & Kamel, 1999). Neural networks with modular designs, termed Modular Neural Networks (MNN), have exhibited a wide range of advantages, including adaptation (Shen et al., 2023b), out-of-distribution (OOD) generalization abilities (Goyal & Bengio, 2020; Weiss et al., 2022), and data efficiency (Bengio et al., 2020). Typical MNNs are usually explicitly modular. Their modular structure is pre-defined and they are expected to achieve a divide-and-conquer solution for the given task. For example, Andreas et al. (2016) design separate visual and language functional modules and jointly train them to solve visual question-answering tasks. Among various MNNs, Mixture-of-Experts (MoEs) employ a conditional computation strategy where different submodules - so-called experts - are expected to be activated by different types of inputs. MoEs sees substantial success in various domains (Shen et al., 2023a; Chen et al., 2023b; Mustafa et al., 2022; Bao et al., 2022) in the era of large-scale transformers, and therefore becomes a widespread modular neural architecture.
@@
Apart from developing explicit MNN,  Csordás et al. (2021); Agarwala et al. (2021) study whether standard neural networks become implicitly modular after training and discover spontaneously emerged modular structure in small-scale CNNs and LSTMs. For more complicated large-scale pre-trained transformers, initial observations (Zhang et al., 2022b; Li et al., 2022) reveal notable sparse activation patterns within the Feed-Forward Networks (FFNs). Specifically, they find that in T5-Base (Raffel et al., 2020) and ViT-B16, only 3.0% and 6.3% neurons are activated during one forward process. However, sparsity does not imply modularity. Sparsely activated FFNs neurons do not ensure to have remarkable functional division. Therefore,  Zhang et al. (2023) further utilize handpicked semantic and knowledge-intensive tasks to probe the nature of neurons in FFNs. They observe a strong correlation between activation and specific tasks, further discovering clear function-based neuron grouping of the model and summarizing such phenomenon as Emergent Modularity.
Figure 1: (a) Existing literatures (Geva et al., 2021; 2022) suggest that the FFNs in transformers can be viewed as key-value memories. They regarded the input as a query, the first layer as keys, and the second as values. Given an input, keys are sparsely activated (marked in red). Most of the values don’t impact the output. (2) The FFNs block can be partitioned into experts by clustering keys. (3) Afterward, experts’ keys averages are used as the gating weights. The inner product between x and gating weights are used to select experts.
However, contemporary transformers are still employed as monolithic models, with their emergent modularity underutilized. Motivated by the advantages of explicit modularity, we explore whether and how standard pre-trained transformers could benefit from emergent modular structures. Technically, we externalize the emergent modularity into the explicit MoEs, i.e., Emergent Mixcture-of-Experts (EMoE). As shown in Figure 1, we first (1) utilize a clustering-based method to decompose the FFNs into experts and then (2) use the Average-Key-Gating ($\operatorname{avg-k}$ gating) to route the input to corresponding experts. Both two steps do not require extra training processes and data.
Then we fine-tune EMoE on a wide range of downstream tasks to investigate if standard pre-trained transformers can benefit from its ermgent modular structure. We evaluate EMoE following different configurations: (1) tasks from various modalities: vision (Domainbed (Gulrajani & Lopez-Paz, 2021)) and language (GLUE (Wang et al., 2019) and GLUE-X (Yang et al., 2023)); (2) different pre-trained dense transformers: ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), and GPT2 (Radford et al., 2019), ranging from 22M to 1.5B parameters. (3) different evaluation settings: in-domain (ID) evaluation and OOD evaluation; (4) different fine-tuning methods: full fine-tuning and parameter-efficient tuning (Hu et al., 2022). In total, we fine-tune 1785 models.
Refer to caption Refer to caption
Figure 2: EMoE achieves stronger ID and OOD performances compared with baselines in both full-fine-tuning (Left, average accuracy) and LoRA tuning settings (Hu et al., 2022) (Right, average accuracy for ID and Friedman rank (Friedman, 1940) for OOD). Results are summarized from Section 4.
Our results are summarized in Figure 2. We find that fine-tuning EMoE achieves significant improvement compared with the vanilla fine-tuning and competitive performance with strong baseline GMoE (Li et al., 2023), demonstrating that dense pre-trained transformer can indeed benefit from emergent modular structures to achieve better generalization performance. We further find that EMoE can deactivate neurons with negative transfer effects, thus improving downstream performance. Meanwhile, our ablation studies indicate that EMoE’s insensitivity to various hyper-parameter configurations and higher data efficiency of up to 20%.
2 Related Work
Modular neural networks has various advantages, including interpretability (Pfeiffer et al., 2023), scalability (Chowdhery et al., 2022), multi-task learning abilities (Chen et al., 2023a), and OOD generalization abilities (Goyal et al., 2021; Li et al., 2023). MoEs (Szymanski & Lemmon, 1993) is currently regarded as a standard framework for developing modular neural networks (Shen et al., 2023b). Most MoEs models are developed by training explicit MoEs architecture from random intialization (Shazeer et al., 2017; Fedus et al., 2022b; Zhang et al., 2022a), which usually causes load imbalance issues (Zoph et al., 2022; Fedus et al., 2022a; Chi et al., 2022). Some recent works employ already-trained dense models as a good initialization (Nie et al., 2021). For example, GMoE (Li et al., 2023) and Upcycling Komatsuzaki et al. (2023) copy the FFNs from a trained dense network to continue pre-training or transfer learning. EMoE is also based on pre-trained dense models. However, unlike existing methods that often significantly increase in parameters (GMoE increases 50% parameters) or tend to be less effective (Upcycling extends T-5 Base (248M) to 2B parameters but still underperforms the T5-Large (783M)), EMoE achieves significant improvements without introducing extra training processes or parameters. This is primarily attributed to the heterogeneity of its experts and gating mechanisms, as we demonstrate in Section  5 and 6.
In another research line,  Hod et al. (2021); Csordás et al. (2021) explore modular structures in CNNs and LSTMs. Some works (Zhang et al., 2022b; Li et al., 2022) focus on the sparsity of more complicated dense pre-trained transformers. Based on their observations,  Zhang et al. (2023) recently explore modularity in pre-trained transformer FFNs. In this work, we adopt clustering-based expert construction and $\operatorname{avg-k}$ gating from MoEfication (Zhang et al., 2022b), while our motivation differs. Seeking to preserve the original functionality with fewer activated neurons, they split every layer in the dense model into MoEs. However, the results are worse than the dense model by more than ten points after splitting (Table 2 in their original paper). This is because the clustering-based decomposition and $\operatorname{avg-k}$ gating cannot precisely select the activated neurons. Without any tuning, these errors accumulate with MoEs layers and lead to severe performance dropping. Therefore, Zhang et al. (2022b) mainly focus on more complicated learning-based construction methods. Differently, we treat these decomposed units as heterogeneous modules to exploit their modular properties in the framework of MoEs, and observe significant improvements over monolithic models.
@@ 3.1 Preliminaries
Transformer Feed-Forward Networks The FFNs layer in the transformer block typically includes weights $\mathbf{K}\in\mathbb{R}^{d\times h}$, $\mathbf{V}\in\mathbb{R}^{h\times d}$, and a non-linear activation function $\sigma(\cdot)$. For an input $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ can be calculated as follows:
$\mathbf{y}=\operatorname{FFN}(\mathbf{x};\mathbf{K},\mathbf{V})=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}.$ (1)
More precisely, for each row $\mathbf{K}_{i,:}$ and column $\mathbf{V_{:,i}}$, the Equation 1 can be rewritten as:
$\mathbf{y}=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}=\sum_{i=1}^{h}\sigma(\mathbf{x}\cdot\mathbf{K_{i,:}})\cdot\mathbf{V_{:,i}}$ (2)
Regarding each row in $\mathbf{K}$ as a key vector and each column in $\mathbf{V}$ as the value vector, the output of an FFNs network can be viewed as a weighted sum of values based on the activation scores $\sigma(\mathbf{x}\cdot\mathbf{K}^{T})$.
Mixture-of-Experts In transformers, MoEs is often implemented by replacing the original FFNs with a group of parallel FFNs and introducing a gating module. Supposing there are $N$ experts: $\{\operatorname{FFN}_{n}(\cdot;\mathbf{K}^{n},\mathbf{V}^{n})|\,n\in[1,N]\}$, the gating module $g(\cdot\,;\mathbf{G},k)$, defined with its parameters $\mathbf{G}$ and an integer $k$, is to map input $\mathbf{x}$ to a score distribution of experts $g(\mathbf{x};\,\mathbf{G},k)\in\mathcal{R}^{N}$. Typically, $g$ is implemented with a simple linear layer followed by a $\operatorname{softmax}$ function and a $\operatorname{Top-k}$ function. Given an input $\mathbf{x}\in\mathbb{R}^{h}$, let $\mathbf{y}_{n}=\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})$ be the output of the $n$-th expert, and then the output $\mathbf{y}\in\mathbb{R}^{h}$ of can also be summarized as the weighted sum of the output from all experts:
$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)y_{n},$ (3)
When $k$ for $\operatorname{Top-K}$ is smaller than $N$, only a subgroup of experts is involved in the computation.
@@ 3.2 Emergent Mixture-Of-Expert
The ideal EMoE should reflect the implicit modular nature of dense pre-trained transformers using the form of MoEs. A preferred approach to constructing EMoE should avoid introducing additional parameters, training, and data, which may result in undesired biases and are impractical. Although more sophisticated constructing methods can be proposed, we adopt a method from MoEfication Zhang et al. (2022b). The method contains two steps: (1) clustering-based expert construction and (2) $\operatorname{avg-k}$ gating. It is simple and fulfills our needs, as demonstrated in our experiments.
Clustering-based Experts Construction. Since key-value pairs with similar key vectors tend to be co-activated, we split them into separate experts according to their key vectors. Specifically, given a trained FFNs layer $\operatorname{FFNs}(\cdot;\mathbf{K},\mathbf{V})$, we perform constrained clustering (Malinen & Fränti, 2014) to partition all key vectors $\mathbf{K}$ into $N$ experts on average, so each group contains $\frac{d}{N}$ key-value pairs (usually $d=4h$). Denoting the indices of keys in the $i$-th group as $E_{i}\subset[d]$, for $\forall j\in E_{i}$, we extract key-value pair $(\mathbf{K_{j,:}},\mathbf{V_{:,j}})$ to form the $i$th expert $\operatorname{FFN}(\cdot;\mathbf{K}_{i},\mathbf{V}_{i})$, as depicted in Figure 1(b). After that, the computation of each expert proceeds as Equation 1.
Avg-k Gating. Given an input $\mathbf{x}$ and $N$ experts, a qualified gate for EMoE should route the input $\mathbf{x}$ to the experts who contribute most to the model’s output. We construct the gating module by averaging each expert’s keys. The gating function is usually implemented by a single layer $\mathbf{G}\in\mathcal{R}^{h\times N}$, in the $\operatorname{avg-k}$ gating’s case, the weights in $n$th column $\mathbf{G}_{,:n}$ can be calculated as follows:
@@
And then the corresponding gating score for $i$-th expert is:
$g_{i}(\mathbf{x};\mathbf{G},k)=\begin{cases}&1~{}~{}\text{ if }i\in\operatorname{Top-K}(\mathbf{x}\cdot\mathbf{G}^{T};k,\text{dim=1})\\ &0~{}~{}\text{ else}\end{cases}$ (5)
where $\operatorname{Top-K}(\cdot;k,\text{dim})$ return indices of $k$ largest element of the given input along a given dimension. Using 0,1 score, $\operatorname{avg-k}$ gating reduces the weights’ influence in Equation 3). Notably, observe:
$\mathbf{x}\cdot\mathbf{G}_{:,n}^{T}=\mathbf{x}\cdot\operatorname{Avg}(\mathbf{K}^{n},\text{dim}=0)=\mathbf{x}\cdot\frac{N}{h}\sum_{j\in E_{i}}\mathbf{K}^{n}_{j,:}=\frac{N}{h}\sum_{j}a_{j}.$ (6)
A larger value of gating score $g_{i}$ implies more activated keys within the corresponding expert. Consequently, the expert could potentially contribute more to the output $\mathbf{y}$ for input $\mathbf{x}$. During downstream tuning, gating weights are updated along with the FFNs parameters using Equation 4.
3.3 Discussion
Though adopting a method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios. MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons. However, the method introduces deviations from FFNs as it can mistakenly eliminate useful neurons. As these deviations accumulated across layers, the performances severely drop by over 10 points on original tasks. Therefore, they focus on more precise expert construction and gating to preserve performances. In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks. We summarize their main differences in Appendix A.1. As one of our most related work, GMoE (Li et al., 2023) introduce MoEs by copying the trained FFNs. They further demonstrate introducing MoEs every 2 layers leads to significant degradation. Accordingly, we only introduce EMoEs in certain layers since efficiency isn’t our primary concern. We also conduct ablation studies about MoEfying every 2 layers and get consistent results, more details are in Appendix A.4.2 Table 14.

### v2

# Unlocking Emergent Modularity in Large Language Models

@@ Abstract
Modular Neural Networks (MNNs) demonstrate various advantages over monolithic models. Existing MNNs are generally explicit: their modular architectures are pre-defined, with individual modules expected to implement distinct functions. Recent works reveal that there exists implicit modularity in standard pre-trained transformers, namely Emergent Modularity. They indicate that such modular structures spontaneously exhibit during the early pre-training phase. Despite the benefits of modularity, most Language Models (LMs) are still treated as monolithic models in the pre-train and fine-tune paradigm, with their emergent modularity locked and underutilized. In this work, focusing on unlocking the emergent modularity in LMs, we showcase that standard LMs could be fine-tuned as their Mixture-of-Expert (MoEs) counterparts without introducing any extra parameters. Such MoEs are derived from emergent modularity and are referred to as Emergent MoEs (EMoE). Our experiments demonstrate that fine-tuning EMoE effectively improves downstream in-domain and out-of-domain generalization compared with vanilla fine-tuning. Our analysis and ablation studies further illustrate that it is robust to various configurations and can scale up to Large Language Models (i.e., Llama2-7B and Llama-30B). Code is available at this repo.
Unlocking Emergent Modularity in Large Language Models
12Zihan Qiuthanks: Equal contributionthanks: Work done while interning at HKUST. , 3Zeyu Huang, 1Jie Futhanks: Corresponding author 1CSE, HKUST     2IIIS, Tsinghua University, 3ILCC, University of Edinburgh qzh11628@gmail.com, zeyu.huang@ed.ac.uk, jiefu@ust.hk
@@ 1 Introduction
Modularity attracts considerable attention from the artificial intelligence community (Auda and Kamel, 1999). Neural networks with modular designs, termed Modular Neural Networks (MNNs), have exhibited a wide range of advantages, including adaptation (Shen et al., 2023b), data efficiency (Bengio et al., 2020), and better generalization abilities (Goyal and Bengio, 2020; Weiss et al., 2022). Typical MNNs are usually explicitly modular. They have a pre-defined modular structure and are expected to achieve a divide-and-conquer solution for the given task. Among various MNNs, Mixture-of-Experts (MoEs) employ a conditional computation strategy where different submodules - so-called experts - are expected to be activated by different types of inputs. MoEs see substantial success in various domains (Fedus et al., 2022; Shen et al., 2023a; Chen et al., 2023b; Mustafa et al., 2022; Bao et al., 2022) in the era of large-scale transformers, and they are therefore a widespread modular neural architecture.
@@
Apart from explicit MNNs, some research finds that modular structures spontaneously emerge during training, not only in small-scale CNNs or LSTMs (Csordás et al., 2021; Agarwala et al., 2021), but also large-scale pre-trained transformer models. Zhang et al. (2022b); Li et al. (2022) reveal notable sparse activation patterns within the Feed-Forward Networks (FFNs) in pre-trained transformer models. They find that in T5-Base (Raffel et al., 2020) and ViT-B16, only 3.0% and 6.3% neurons are activated during one forward process, respectively. Furthermore,  Zhang et al. (2023) utilize handpicked semantic and knowledge-intensive tasks to probe the nature of neurons in FFNs. They observe a strong correlation between neuron activation and specific tasks, further discovering clear function-based neuron grouping of the pre-trained T5 model (neurons with similar functions are usually co-activated). They summarize such phenomenon as Emergent Modularity (EM).
Though modularity emerges, pre-trained language models are generally treated as monolithic models in the standard pre-train and fine-tune paradigm. It is natural to ask whether their EM and the potential improvements brought by EM are locked in this process.
In this paper, we advocate unlocking the EM in the pre-trained language models could bring generalization improvements for downstream tasks. Specifically, we split certain FFNs layers of the original model into MoEs layers. The MoEs is derived according to the EM in that layer and can be regarded as the externalization of the EM. Hence, the obtained MoEs model is called Emergent MoEs (EMoE). We then fine-tune the obtained EMoE model to investigate whether unlocking EM encourages downstream task performance.
We validate our empirical findings with various models, evaluation benchmarks, fine-tuning methods (parameter-efficient tuning and full fine-tuning). We find that fine-tuning EMoE achieves stronger generalization performance than vanilla fine-tuning across various experimental settings, demonstrating that unlocking the EM of LMs boosts the models’ downstream generalization abilities. We provide a comprehensive analysis for EMoE: 1) We first validate that EMoE indeed unlocks EM in pre-trained language models by showcasing its task-specific expert choice. 2) We then reveal that EMoE ameliorates the parameter updating during fine-tuning and can even be abandoned afterward. We want to highlight that this property further improves the practicality of EMoE as the model architecture does not change before and after fine-tuning. Meanwhile, our ablation studies show the EMoE’s robustness to various hyper-parameter configurations. 3) We also conclude that EMoE could mask neurons with negative transfer effects. We hope our research discoveries can bring novel insights and serve as an example attempt towards further unlocking the EM of LLMs.
@@ 2.1 Preliminaries
Transformer FFNs are Key-Value Memories. The FFNs layer in the transformer block typically includes weights $\mathbf{K}\in\mathbb{R}^{h\times d}$, $\mathbf{V}\in\mathbb{R}^{d\times h}$, where $h$ is embedding size and $d$ is the dimension of the hidden layer (usually $d=4h$), and a non-linear activation function $\sigma(\cdot)$. For an input $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ can be calculated as Eq. 1:
$\mathbf{y}=\operatorname{FFN}(\mathbf{x};\mathbf{K},\mathbf{V})=\sigma(\mathbf{x}\cdot\mathbf{K})\cdot\mathbf{V}.$ (1)
More precisely, for each column $\mathbf{K}_{:,i}$ and row $\mathbf{V}_{i,:}$, Eq.1 can be rewritten as:
$\mathbf{y}=\sigma(\mathbf{x}\cdot\mathbf{K})\cdot\mathbf{V}=\sum_{i=1}^{h}\sigma(\mathbf{x}\cdot\mathbf{K}_{:,i})\cdot\mathbf{V}_{i,:}$ (2)
Following Geva et al. (2021, 2022); Huang et al. (2023), we regard columns in $\mathbf{K}$ as key vectors and rows in $\mathbf{V}$ as value vectors, the output of an FFNs network can be viewed as a weighted sum of value vectors based on the activation scores $\sigma(\mathbf{x}\cdot\mathbf{K})$. For the rest of this paper, we refer to one key-value memory pair using neuron and the co-activation of neurons using modularity.
Mixture-of-Experts In transformers, MoEs is often implemented by replacing the original FFNs with a group of parallel FFNs and introducing a gating module. Supposing there are $N$ experts: $\{\operatorname{FFN}^{n}(\cdot;\mathbf{K}^{n},\mathbf{V}^{n})|\,n\in[1,N]\}$, the gating module $g(\cdot\,;\mathbf{G},k)$, defined with its parameters $\mathbf{G}$ and an integer $k$, is to map input $\mathbf{x}$ to a score distribution of experts $g(\mathbf{x};\,\mathbf{G},k)\in\mathcal{R}^{N}$. Typically, $g$ is implemented with a simple linear layer followed by a $\operatorname{softmax}$ function and a $\operatorname{Top-k}$ function. Given $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ of can be summarized as the weighted sum of the output from all experts:
$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)\operatorname{FFN}^{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})$ (3)
When $k$ for $\operatorname{Top-K}$ is smaller than $N$, only a subgroup of experts is involved in the computation, termed sparse MoEs.
@@ 2.2 Emergent Mixture-of-Expert
According to Eq. 2 and Eq. 3, we find that FFNs internally resemble MoEs if we consider keys as the gating module and values as the group of experts, which inspire us to transform existing FFNs into sparse MoEs to unlock its modular potentials. Since our research goals mainly focus on EM, a preferred approach to externalizing EM into sparse MoEs should not introduce additional parameters, training, and data, which may result in impractical or undesired biases. Therefore, after splitting original neurons into different groups to construct different experts, each group’s average of key vectors is calculated to form the gating module. Details are described and illustrated in Fig. 1.
Figure 1: (a) Existing literatures (Geva et al., 2021, 2022) suggest that the FFNs in transformers can be viewed as key-value memories. They regarded the input as a query, the first layer as keys, and the second as values. Given an input, keys are sparsely activated (marked in red). Most of the values don’t impact the output. (2) The FFNs block can be partitioned into experts by clustering keys. (3) Afterward, experts’ key averages are used as the gating weights. The inner product between x and gating weights are used to select experts.
Clustering-based Experts Construction. We aim to ensure that neurons that tend to be co-activated are divided into the same group. Since neurons with similar key vectors tend to be co-activated according to Eq. 2, we split them into separate experts by clustering their key vectors. Specifically, given a FFNs layer $\operatorname{FFN}(\cdot;\mathbf{K},\mathbf{V})$, we perform constrained clustering (Malinen and Fränti, 2014) (more details are in Appendix A.0.1) to partition all key vectors $\mathbf{K}$ into $N$ experts on average. Denoting the indices of keys in the $i$-th group as $E_{i}\subset[d]$, for $\forall j\in E_{i}$, we extract neuron $(\mathbf{K}_{:,j},\mathbf{V}_{j,:})$ to form the $i$-th expert $\operatorname{FFN}(\cdot;\mathbf{K}^{i},\mathbf{V}^{i})$ as depicted in Fig. 1(b). After that, the computation of each expert follows Eq. 1.
Avg-k Gating. As discussed, we do not want to introduce extra trainable parameters when externalizing EM. Therefore, we construct the gating module by averaging each expert’s keys, which should route the input $\mathbf{x}$ to the experts who tend to have larger activation scores and thus contribute more to the model’s output. The gating function is usually implemented by a single layer $\mathbf{G}\in\mathcal{R}^{h\times N}$, in the $\operatorname{avg-k}$ gating’s case, the weights in $i$-th column $\mathbf{G}_{:,i}$ can be calculated as follows:
@@
And then the gating score for $i$-th expert is:
$g_{i}(\mathbf{x};\mathbf{G},k)=\begin{cases}&1~{}~{}\text{ if }i\in\operatorname{Top-K}(\mathbf{x}\cdot\mathbf{G};k)\\ &0~{}~{}\text{ else}\end{cases}$ (5)
where $\operatorname{Top-K}(\cdot;k)$ returns indices of $k$ largest element of the given input along. As the gating score is the average of activation scores (before $\sigma(\cdot)$) of neurons in that expert:
$\mathbf{x}\cdot\mathbf{G}_{:,i}=\mathbf{x}\cdot\operatorname{Avg}(\mathbf{K}^{i})=\frac{N}{d}\sum_{j\in E_{i}}\mathbf{x}\mathbf{K}^{i}_{:,j}=\frac{N}{d}\sum_{j}a_{j}.$ (6)
a larger value of gating score $g_{i}$ implies more activated keys within the corresponding expert. Consequently, the expert could potentially contribute more to the output $\mathbf{y}$ for input $\mathbf{x}$. During downstream tuning, gating weights are tied with the FFNs parameters using Eq. 4.